You receive a prescription for a drug to be given in a dose of 5 mg/m2. Calculate the dose of the drug for a patient with a BSA of 1.25m2. Solve the problem using both ratio and proportion and dimensional analysis.
1
Expert's answer
2012-09-28T09:05:05-0400
a) using ratio and proportion analysis:
5 / 1 = x / 1.25
x = 5 * 1.25 =
6.25 mg
b) using dimensional analysis: [x] = [ (mg/m2) * m2 ] = [mg]